OpenStudy (kaley246):

Graph the two equations to solve the system. What is the ordered pair solution? x + y = 6 and x - y = 4 (Points : 1) (4, 2) (6, 2) (5, 1) no solution

OpenStudy (kaley246):

@robtobey

OpenStudy (mayankdevnani):

there are two methods :- 1) to solve it by graphical method ( long method) 2) by algebraically ( short method )

OpenStudy (kaley246):

short method

OpenStudy (mayankdevnani):

\[\large \bf x-y=4\] \[\large \bf x=4+y\]

OpenStudy (mayankdevnani):

put this value of x and plug in the first equation !

OpenStudy (mayankdevnani):

we do this because we need to solve two equations and to find ordered pair ( common) , we use this ^^

OpenStudy (mayankdevnani):

understood ? @kaley246

OpenStudy (kaley246):

no i like how robtobey does it

OpenStudy (mayankdevnani):

as you wish !

OpenStudy (kaley246):

your way is too complianed

OpenStudy (kaley246):

i am sorry but i am special i have promblem understanding dtuff

OpenStudy (mayankdevnani):

its depend upon user to user !

OpenStudy (mayankdevnani):

carry on ! @kaley246

OpenStudy (kaley246):

ok robtobey help please

OpenStudy (kaley246):

ok @robtobey help please

OpenStudy (anonymous):

Sorry. The first posting had an error. x+y=6 x-y=4 --------- 2 x = 10 x = 5

OpenStudy (anonymous):

To find y: 5 + ? = 6 5 + y = 6 y = 1